# Towards Understanding Why FixMatch Generalizes Better Than Supervised Learning

### v1

@@ E.1 Useful Claims
$\quad-O\left(\frac{\eta\tilde{N}_{u,s}}{\tilde{N}_{u}}\right)\mathbb{E}_{(X,y)\sim\tilde{\mathcal{Z}}_{u,s}}\big{[}\tilde{O}(\sigma_{p}P)\mathbb{I}_{y=i}(1-\mathbf{logit}_{i}(F^{(t)},\mathcal{A}(X)))\big{]}$
$\quad-O\left(\frac{\eta\tilde{N}_{u,s}}{\tilde{N}_{u}}\right)\mathbb{E}_{(X,y)\sim\tilde{\mathcal{Z}}_{u,s}}\big{[}\mathbb{I}_{y\neq i}(\mathcal{E}_{1}+\mathcal{E}_{3}+\mathbb{I}_{v_{i,3-l^{*}}\in\mathcal{P}(X)})\mathbf{logit}_{i}(F^{(t)},\mathcal{A}(X))\big{]}$
$\geq\langle w_{i,r}^{(t)},v_{i,3-l^{*}}\rangle+\Omega\left(\frac{\eta}{k}\right)\mathbb{E}_{(X,y)\sim\tilde{\mathcal{Z}}_{u,m}}\left[(1-\mathbf{logit}_{y}(F^{(t)},\mathcal{A}(X)))\right]-O\left(\frac{\eta s\tilde{N}_{u,s}}{k^{2}\tilde{N}_{u}}\right),$
where the last step is based on Claim 23 and the fact that $\Pr(v_{i,l}\in\mathcal{P}(X)|i\neq y)=\frac{s}{k}\ll o(1)$. Thus, when summing up all $r\in[m]$, and telescoping from $T+T_{0}$ to $T+T^{\prime}$, we have
$\Phi_{i,3-l^{*}}^{(T^{\prime}+T)}\geq\Phi_{i,3-l^{*}}^{(T+T_{0})}+\tilde{\Omega}\left(\frac{\eta}{k}\right)\sum_{t=T+T_{0}}^{T+T^{\prime}}\mathbb{E}_{(X,y)\sim\tilde{\mathcal{Z}}_{u,m}}\left[(1-\mathbf{logit}_{y}(F^{(t)},\mathcal{A}(X)))\right]-\tilde{O}\left(\frac{T^{\prime}\eta s\tilde{N}_{u,s}}{k^{2}\tilde{N}_{u}}\right).$
@@
$\sum_{t=T+T_{0}}^{T+T^{\prime}}\mathbb{E}_{(X,y)\sim\tilde{\mathcal{Z}}_{u,m}}\left[(1-\mathbf{logit}_{i}(F^{(t)},\mathcal{A}(X)))\right]\leq\tilde{O}\left(\frac{k}{\eta}\right)+\tilde{O}\left(\frac{T^{\prime}s\tilde{N}_{u,s}}{k\tilde{N}_{u}}\right)\leq\tilde{O}\left(\frac{k}{\eta}\right).$
@@
Suppose that the Induction Hypothesis 15 holds for every iteration $T<t\leq T+T^{\prime}$, under Parameter Assumption 10 and 12, we have for every iteration $T<t\leq T+T^{\prime}$:
$\forall(i,l^{*})\in\mathcal{M}_{F},\quad\Phi_{i,l^{*}}^{(t)}\geq\Omega(\log(k).$
@@
Recall $V_{i,r,l^{*}}(X):=\mathbb{I}_{v_{i,l^{*}}\in\mathcal{V}(X)}\sum_{p\in\mathcal{P}_{v_{i,l^{*}}}(X)}\overline{\operatorname{ReLU}}^{\prime}(\langle w_{i,r}^{(t)},\mathcal{A}(x_{p})\rangle)\mathcal{A}(z_{p})$ and our definition of strong augmentation in (14), we have $V_{i,r,l^{*}}(X)=0$, so we have by Claim 23 that
@@
In the following, we separate the process into $T\leq t\leq T+T_{0}$ and $t\geq T_{0}+T$.
When $T<t\leq T+T_{0}$.
Recall at the end of learning Phase I, we have $\Phi_{i,l^{*}}^{(T)}\geq\Omega(\log(k)$. Using $T_{0}=\tilde{\Theta}\left(\frac{k}{\eta\sigma_{0}^{q-2}}\right)$ and our parameter assumption 10, we have $\Phi_{i,l^{*}}^{(t)}\geq\Omega(\log(k)$ for every iteration of $T<t\leq T+T_{0}$.
When $T+T_{0}<t\leq T+T^{\prime}$.
By the upper bound on multi-view error in Claim 26, we know $\Phi_{i,l^{*}}^{(t)}\geq\Omega(\log(k)$ for every iteration of $T+T_{0}<t\leq T+T^{\prime}$. ∎
@@
When $\mathbb{E}_{(X,y)\sim\tilde{\mathcal{Z}}_{u,m}}\left[1-\mathbf{logit}_{y}\left(F^{(t)},X\right)\right]\leq\frac{1}{k^{4}}$ is sufficiently small, we have for any $(i,3-l),(j,3-l^{\prime})\notin\mathcal{M}_{F}$,
@@
Denote by $\tilde{\mathcal{Z}}^{*}_{u,m}$ for the sample $(X,y)\in\tilde{\mathcal{Z}}_{u,m}$ such that $\sum_{p\in P_{v_{y,3-l^{*}}}(X)}z_{p}\leq 1+\frac{1}{100\log(k)}$ where $(y,l^{*})\in\mathcal{M}_{F}$. For a sample $(X,y)\in\tilde{\mathcal{Z}}^{*}_{u,m}$, denote by $\mathcal{H}(X)$ as the set of all $i\in[k]\setminus\{y\}$ such that $\sum_{p\in P_{v_{i,3-l}}(X)}z_{p}\geq 0.4-\frac{1}{100\log(k)}$ where $(i,l)\in\mathcal{M}_{F}$.
@@
where $(j,l^{\prime})\in\mathcal{M}_{F}$. Note for every the probability for every $i\neq j\in[k]$, the probability of generating a sample $(X,y)\in\tilde{\mathcal{Z}}^{*}_{u,m}$ with $y=j$ and $i\in\mathcal{H}(X)$ is at least $\tilde{\Omega}(\frac{1}{k}\cdot\frac{s^{2}}{k^{2}})$. This implies
@@
By symmetry and non-negativity of $\Phi_{i,3-l}^{(t)}$, we know for any $(i,3-l),(j,3-l^{\prime})\notin\mathcal{M}_{F}$, we have:
@@
Since (16) holds for any $(i,3-l),(j,3-l^{\prime})\notin\mathcal{M}_{F}$ at iteration $t$ such that $\mathbb{E}_{(X,y)\sim\tilde{\mathcal{Z}}_{m}}\left[1-\mathbf{logit}_{y}\left(F^{(t)},X\right)\right]\leq\frac{1}{k^{4}}$, and from Claim D.16 in Allen-Zhu & Li (2023) we know (16) also holds for $\Phi_{i,l},\Phi_{j,l^{\prime}}$ for any $(i,l),(j,l^{\prime})\in\mathcal{M}_{F}$ during learning Phase II, so we have
@@
Let $T+T_{1}$ be the first iteration that $\mathbb{E}_{(X,y)\sim\tilde{\mathcal{Z}}_{u,m}}\left[1-\mathbf{logit}_{y}\left(F^{(t)},X\right)\right]\leq\frac{1}{k^{4}}$, then we know for $t\geq T+T_{1}$ (16) always holds, since the objective $L^{(t)}=L_{s}^{(t)}+\lambda L_{u}^{(t)}$ ($\lambda=1$) is $O(1)$-Lipschitz smooth and we are using full gradient descent, which means the objective value is monotonically non-increasing. Since in Phase II, $L_{s}^{(t)}$ is kept at a small value, $L_{u}^{(t)}$ is monotonically non-increasing.
@@ E.2.1 Correlation Growth
Suppose Parameter 10 holds and suppose Induction Hypothesis 15 holds for all iteration $<t$ starting from $T$. Then, letting $\Phi_{i,l}^{(t)}:=\sum_{r\in[m]}[\langle w_{i,r}^{(t)},v_{i,l}\rangle]^{+}$, we have for every $i\in[k],l\in[2]$,
@@
For every $i\in[k]$ and every $(i,l)\in\mathcal{M}_{F}$, after the first stage, we have $\Phi_{i,l}^{(T)}\leq\tilde{O}(1)$. In the second stage, as in Semantic-Aware CutOut, the learned features $(i,l)$ are masked and so the correlations between gradients and learned features are kept small. This means that $\Phi_{i,l}^{(t)}\leq\tilde{O}(1)$ holds true in learning Phase II for $T<t\leq T+T^{\prime}$.
Then for the unlearned feature $(i,3-l)$, we suppose $t>T+T_{1}$ is some iteration so that $\Phi_{i,3-l}^{(t)}\geq\mathrm{polylog}(k)$. We will prove that if we continue from iteration $t$ for at most $T^{\prime}$ iterations, we still have $\Phi_{i,3-l}^{(t)}\leq\tilde{O}(1)$. Based on Claim 19, we have that
@@
$\leq\langle w_{i,r}^{(t)},v_{i,3-l}\rangle+O(\eta)\mathbb{E}_{(X,y)\sim\tilde{\mathcal{Z}}_{u,m}}\left[\mathbb{I}_{i=y}(1-\mathbf{logit}_{i}(F^{(t)},\mathcal{A}(X)))\right]+O\left(\frac{\eta\rho\tilde{N}_{u,s}}{k\tilde{N}_{u}}\right).$
@@
Then when $j\neq y$ and $(j,l^{\prime})\in\mathcal{M}_{F}$, we have
@@
$\Phi_{i,3-l}^{(t+1)}\leq\Phi_{i,3-l}^{(t)}+\frac{\eta m}{k^{\Omega(\log k)}}+\tilde{O}\left(\frac{\eta\rho\tilde{N}_{u,s}}{k\tilde{N}_{u}}\right).$
Therefore, if we continue for $T^{\prime}$ iterations, we still have $\Phi_{i,3-l^{*}}^{(T+T^{\prime})}\leq\tilde{O}(1)$. ∎
@@ E.2.2 Off-Diagonal Correlations are Small
Suppose Parameter 10 holds and suppose Induction Hypothesis 15 holds for all iteration $<t$ starting from $T$. Then,
@@
In Phase I when $t\leq T$, from Lemma D.22 in Allen-Zhu & Li (2023), we have $|\langle w_{i,r}^{(t)},v_{j,l}\rangle|\leq\tilde{O}(\sigma_{0})$. Now we consider Phase II when $t>T$, and denote by $R_{i}^{(t)}:=\max_{r\in[m],j\in[k]\setminus\{i\}}|\langle w_{i,r}^{(t)},v_{j,l}\rangle|$. According to Claim 19 and Claim 20, we have
@@
For single-view data $(X,y)\sim\tilde{\mathcal{Z}}_{u,s}$, by Claim 23, we have $\mathbf{logit}_{i}(F^{(t)},\mathcal{A}(X))=O\left(\frac{1}{k}\right)$ for every $i\in[k]$. In the following, we separate the process into $T<t\leq T+T_{0}$ and $T+T_{0}<t\leq T+T^{\prime}$.
When $T<t\leq T+T_{0}$.
During this stage, by Claim 23 we know $\mathbf{logit}_{i}(F^{(t)},X)=O(\frac{1}{k})$ ($\forall i\in[k]$) for any $(X,y)\sim\tilde{\mathcal{Z}}_{u}$. We also have $\mathcal{E}_{2,i,r}(X)\leq\tilde{O}(\gamma(\Phi_{i,3-l^{*}}^{(t)})^{q-1})\text{ with }(i,l^{*})\in\mathcal{M}_{F},\text{ and have }\mathcal{E}_{4,j,l}(X)\leq\tilde{O}(\sigma_{0})^{q-1}\mathbb{I}_{v_{j,l}\in\mathcal{V}(X)}$ by definition. Recall when $T<t\leq T+T_{0}$, by Claim 24, we have $\Phi_{i,3-l^{*}}^{(t+1)}=\Phi_{i,3-l^{*}}^{(t)}+\tilde{\Theta}\left(\frac{\eta}{k}\right)\overline{\operatorname{ReLU}}^{\prime}(\Phi_{i,3-l^{*}}^{(t)})$, so $\sum_{T<t\leq T+T_{0}}\eta(\Phi_{i,3-l^{*}}^{(t)})^{q-1}\leq\tilde{O}(k)$. Also, $\Pr(v_{i,3-l^{*}}\in\mathcal{P}(X)|i\neq y)=\frac{s}{k}$. Therefore, for every $T<t\leq T+T_{0}$ with $T_{0}=\tilde{\Theta}(\frac{k}{\eta\sigma_{0}^{q-2}})$, we have
$R_{i}^{(t)}\leq R_{i}^{(T)}+\tilde{O}(\sigma_{0})+\tilde{O}\left(\frac{\eta}{k}T_{0}\right)\left((\sigma_{0}^{q-1})\gamma s+(\sigma_{0}\gamma k)^{q-1}\gamma P+(\sigma_{0})^{q-1}\frac{s}{k}\right)\leq\tilde{O}(\sigma_{0}).$
When $T+T_{0}<t\leq T+T^{\prime}$.
@@
Therefore, by the upper bound on multi-view error in Claim 26 and $\frac{\tilde{N}_{u,s}}{\tilde{N}_{u}}\leq\frac{k^{2}}{\eta sT^{\prime}}$, we know $R_{i}^{(t)}\leq\tilde{O}(\sigma_{0})$ for $T+T_{0}<t\leq T+T^{\prime}$. ∎

### v2

@@ E.1 Useful Claims
$\quad-O\left(\frac{\eta\tilde{N}_{c,s}}{\tilde{N}_{c}}\right)\mathbb{E}_{(X,y)\sim\tilde{\mathcal{Z}}_{u,s}}\big{[}\tilde{O}(\sigma_{p}P)\mathbb{I}_{y=i}(1-\mathbf{logit}_{i}(F^{(t)},\mathcal{A}(X)))\big{]}$
$\quad-O\left(\frac{\eta\tilde{N}_{c,s}}{\tilde{N}_{c}}\right)\mathbb{E}_{(X,y)\sim\tilde{\mathcal{Z}}_{u,s}}\big{[}\mathbb{I}_{y\neq i}(\mathcal{E}_{1}+\mathcal{E}_{3}+\mathbb{I}_{v_{i,3-l^{*}}\in\mathcal{P}(X)})\mathbf{logit}_{i}(F^{(t)},\mathcal{A}(X))\big{]}$
$\geq\langle w_{i,r}^{(t)},v_{i,3-l^{*}}\rangle+\Omega\left(\frac{\eta}{k}\right)\mathbb{E}_{(X,y)\sim\tilde{\mathcal{Z}}_{u,m}}\left[(1-\mathbf{logit}_{y}(F^{(t)},\mathcal{A}(X)))\right]-O\left(\frac{\eta s\tilde{N}_{c,s}}{k^{2}\tilde{N}_{c}}\right),$
where the last step is based on Claim 23 and the fact that $\Pr(v_{i,l}\in\mathcal{P}(X)|i\neq y)=\frac{s}{k}\ll o(1)$. Thus, when summing up all $r\in[m]$, and telescoping from $T_{1}+T_{0}$ to $T_{1}+T_{2}$, we have
$\Phi_{i,3-l^{*}}^{(T_{1}+T_{2})}\geq\Phi_{i,3-l^{*}}^{(T_{1}+T_{0})}+\tilde{\Omega}\left(\frac{\eta}{k}\right)\sum_{t=T_{1}+T_{0}}^{T_{1}+T_{2}}\mathbb{E}_{(X,y)\sim\tilde{\mathcal{Z}}_{u,m}}\left[(1-\mathbf{logit}_{y}(F^{(t)},\mathcal{A}(X)))\right]-\tilde{O}\left(\frac{T_{2}\eta s\tilde{N}_{c,s}}{k^{2}\tilde{N}_{c}}\right).$
@@
$\sum_{t=T_{1}+T_{0}}^{T_{1}+T_{2}}\mathbb{E}_{(X,y)\sim\tilde{\mathcal{Z}}_{u,m}}\left[(1-\mathbf{logit}_{i}(F^{(t)},\mathcal{A}(X)))\right]\leq\tilde{O}\left(\frac{k}{\eta}\right)+\tilde{O}\left(\frac{T_{2}s\tilde{N}_{c,s}}{k\tilde{N}_{c}}\right)\leq\tilde{O}\left(\frac{k}{\eta}\right).$
@@
Suppose that the Induction Hypothesis 15 holds for every iteration $T_{1}<t\leq T_{1}+T_{2}$, under Parameter Assumption 10 and 12, we have for every iteration $T_{1}<t\leq T_{1}+T_{2}$:
$\forall(i,l^{*})\in\mathcal{M},\quad\Phi_{i,l^{*}}^{(t)}\geq\Omega(\log(k)).$
@@
Recall $V_{i,r,l^{*}}(X):=\mathbb{I}_{v_{i,l^{*}}\in\mathcal{V}(X)}\sum_{p\in\mathcal{P}_{v_{i,l^{*}}}(X)}\overline{\operatorname{ReLU}}^{\prime}(\langle w_{i,r}^{(t)},\mathcal{A}(x_{p})\rangle)\mathcal{A}(z_{p})$ and our definition of strong augmentation in Eq. (14), we have $V_{i,r,l^{*}}(X)=0$, so we have by Claim 23 that
@@
In the following, we separate the process into $T_{1}<t\leq T_{1}+T_{0}$ and $t\geq T_{1}+T_{0}$.
When $T_{1}<t\leq T_{1}+T_{0}$.
Recall at the end of learning Phase I, we have $\Phi_{i,l^{*}}^{(T_{1})}\geq\Omega(\log(k))$. Using $T_{0}=\tilde{\Theta}\left(\frac{k}{\eta\sigma_{0}^{q-2}}\right)$ and our 10, we have $\Phi_{i,l^{*}}^{(t)}\geq\Omega(\log(k))$ for every iteration of $T_{1}<t\leq T_{1}+T_{0}$.
When $T_{1}+T_{0}<t\leq T_{1}+T_{2}$.
By the upper bound on multi-view error in Claim 26, we know $\Phi_{i,l^{*}}^{(t)}\geq\Omega(\log(k))$ for every iteration of $T_{1}+T_{0}<t\leq T_{1}+T_{2}$. ∎
@@
When $\mathbb{E}_{(X,y)\sim\tilde{\mathcal{Z}}_{u,m}}\left[1-\mathbf{logit}_{y}\left(F^{(t)},X\right)\right]\leq\frac{1}{k^{4}}$ is sufficiently small, we have for any $(i,3-l),(j,3-l^{\prime})\notin\mathcal{M}$,
@@
Denote by $\tilde{\mathcal{Z}}^{*}_{u,m}$ for the sample $(X,y)\in\tilde{\mathcal{Z}}_{u,m}$ such that $\sum_{p\in P_{v_{y,3-l^{*}}}(X)}z_{p}\leq 1+\frac{1}{100\log(k)}$ where $(y,l^{*})\in\mathcal{M}$. For a sample $(X,y)\in\tilde{\mathcal{Z}}^{*}_{u,m}$, denote by $\mathcal{H}(X)$ as the set of all $i\in[k]\setminus\{y\}$ such that $\sum_{p\in P_{v_{i,3-l}}(X)}z_{p}\geq 0.4-\frac{1}{100\log(k)}$ where $(i,l)\in\mathcal{M}$.
@@
where $(j,l^{\prime})\in\mathcal{M}$. Note for every the probability for every $i\neq j\in[k]$, the probability of generating a sample $(X,y)\in\tilde{\mathcal{Z}}^{*}_{u,m}$ with $y=j$ and $i\in\mathcal{H}(X)$ is at least $\tilde{\Omega}(\frac{1}{k}\cdot\frac{s^{2}}{k^{2}})$. This implies
@@
By symmetry and non-negativity of $\Phi_{i,3-l}^{(t)}$, we know for any $(i,3-l),(j,3-l^{\prime})\notin\mathcal{M}$, we have:
@@
Since Eq. (16) holds for any $(i,3-l),(j,3-l^{\prime})\notin\mathcal{M}$ at iteration $t$ such that $\mathbb{E}_{(X,y)\sim\tilde{\mathcal{Z}}_{m}}\left[1-\mathbf{logit}_{y}\left(F^{(t)},X\right)\right]\leq\frac{1}{k^{4}}$, and from Claim D.16 in Allen-Zhu & Li (2023) we know Eq. (16) also holds for $\Phi_{i,l},\Phi_{j,l^{\prime}}$ for any $(i,l),(j,l^{\prime})\in\mathcal{M}$ during learning Phase II, so we have
@@
Let $T_{1}+T_{0}^{\prime}$ be the first iteration that $\mathbb{E}_{(X,y)\sim\tilde{\mathcal{Z}}_{u,m}}\left[1-\mathbf{logit}_{y}\left(F^{(t)},X\right)\right]\leq\frac{1}{k^{4}}$, then we know for $t\geq T_{1}+T_{0}^{\prime}$ Eq. (16) always holds, since the objective $L^{(t)}=L_{s}^{(t)}+\lambda L_{u}^{(t)}$ ($\lambda=1$) is $O(1)$-Lipschitz smooth and we are using full gradient descent, which means the objective value is monotonically non-increasing. Since in Phase II, $L_{s}^{(t)}$ is kept at a small value, $L_{u}^{(t)}$ is monotonically non-increasing.
@@ E.2.1 Correlation Growth
Suppose Parameter 10 holds and suppose Induction Hypothesis 15 holds for all iteration $<t$ starting from $T_{1}$. Then, letting $\Phi_{i,l}^{(t)}:=\sum_{r\in[m]}[\langle w_{i,r}^{(t)},v_{i,l}\rangle]^{+}$, we have for every $i\in[k],l\in[2]$,
@@
For every $i\in[k]$ and every $(i,l)\in\mathcal{M}$, after the Phase I, we have $\Phi_{i,l}^{(T_{1})}\leq\tilde{O}(1)$. In Phase II, as in SA-CutOut, the learned features $(i,l)$ are removed and so the correlations between gradients and learned features are kept small. This means that $\Phi_{i,l}^{(t)}\leq\tilde{O}(1)$ holds true in learning Phase II for $T_{1}<t\leq T_{1}+T_{2}$.
Then for the unlearned feature $(i,3-l)$, we suppose $t>T_{1}+T_{0}^{\prime}$ is some iteration so that $\Phi_{i,3-l}^{(t)}\geq\mathrm{polylog}(k)$. We will prove that if we continue from iteration $t$ for at most $T_{2}$ iterations, we still have $\Phi_{i,3-l}^{(t)}\leq\tilde{O}(1)$. Based on Claim 19, we have that
@@
$\leq\langle w_{i,r}^{(t)},v_{i,3-l}\rangle+O(\eta)\mathbb{E}_{(X,y)\sim\tilde{\mathcal{Z}}_{u,m}}\left[\mathbb{I}_{i=y}(1-\mathbf{logit}_{i}(F^{(t)},\mathcal{A}(X)))\right]+O\left(\frac{\eta\rho\tilde{N}_{c,s}}{k\tilde{N}_{c}}\right).$
@@
Then when $j\neq y$ and $(j,l^{\prime})\in\mathcal{M}$, we have
@@
$\Phi_{i,3-l}^{(t+1)}\leq\Phi_{i,3-l}^{(t)}+\frac{\eta m}{k^{\Omega(\log k)}}+\tilde{O}\left(\frac{\eta\rho\tilde{N}_{c,s}}{k\tilde{N}_{c}}\right).$
Therefore, if we continue for $T_{2}$ iterations, we still have $\Phi_{i,3-l^{*}}^{(T_{1}+T_{2})}\leq\tilde{O}(1)$. ∎
@@ E.2.2 Off-Diagonal Correlations are Small
Suppose Parameter 10 holds and suppose Induction Hypothesis 15 holds for all iteration $<t$ starting from $T_{1}$. Then,
@@
In Phase I when $t\leq T_{1}$, from Lemma D.22 in Allen-Zhu & Li (2023), we have $|\langle w_{i,r}^{(t)},v_{j,l}\rangle|\leq\tilde{O}(\sigma_{0})$. Now we consider Phase II when $t>T_{1}$, and denote by $R_{i}^{(t)}:=\max_{r\in[m],j\in[k]\setminus\{i\}}|\langle w_{i,r}^{(t)},v_{j,l}\rangle|$. According to Claim 19 and Claim 20, we have
@@
For single-view data $(X,y)\sim\tilde{\mathcal{Z}}_{u,s}$, by Claim 23, we have $\mathbf{logit}_{i}(F^{(t)},\mathcal{A}(X))=O\left(\frac{1}{k}\right)$ for every $i\in[k]$. In the following, we separate the process into $T_{1}<t\leq T_{1}+T_{0}$ and $T_{1}+T_{0}<t\leq T_{1}+T_{2}$.
When $T_{1}<t\leq T_{1}+T_{0}$.
During this stage, by Claim 23 we know $\mathbf{logit}_{i}(F^{(t)},X)=O(\frac{1}{k})$ ($\forall i\in[k]$) for any $(X,y)\sim\tilde{\mathcal{Z}}_{u}$. We also have $\mathcal{E}_{2,i,r}(X)\leq\tilde{O}(\gamma(\Phi_{i,3-l^{*}}^{(t)})^{q-1})\text{ with }(i,l^{*})\in\mathcal{M},\text{ and have }\mathcal{E}_{4,j,l}(X)\leq\tilde{O}(\sigma_{0})^{q-1}\mathbb{I}_{v_{j,l}\in\mathcal{V}(X)}$ by definition. Recall when $T_{1}<t\leq T_{1}+T_{0}$, by Claim 24, we have $\Phi_{i,3-l^{*}}^{(t+1)}=\Phi_{i,3-l^{*}}^{(t)}+\tilde{\Theta}\left(\frac{\eta}{k}\right)\overline{\operatorname{ReLU}}^{\prime}(\Phi_{i,3-l^{*}}^{(t)})$, so $\sum_{T_{1}<t\leq T_{1}+T_{0}}\eta(\Phi_{i,3-l^{*}}^{(t)})^{q-1}\leq\tilde{O}(k)$. Also, $\Pr(v_{i,3-l^{*}}\in\mathcal{P}(X)|i\neq y)=\frac{s}{k}$. Therefore, for every $T_{1}<t\leq T_{1}+T_{0}$ with $T_{0}=\tilde{\Theta}(\frac{k}{\eta\sigma_{0}^{q-2}})$, we have
$R_{i}^{(t)}\leq R_{i}^{(T_{1})}+\tilde{O}(\sigma_{0})+\tilde{O}\left(\frac{\eta}{k}T_{0}\right)\left((\sigma_{0}^{q-1})\gamma s+(\sigma_{0}\gamma k)^{q-1}\gamma P+(\sigma_{0})^{q-1}\frac{s}{k}\right)\leq\tilde{O}(\sigma_{0}).$
When $T_{1}+T_{0}<t\leq T_{1}+T_{2}$.
@@
Therefore, by the upper bound on multi-view error in Claim 26 and $\frac{\tilde{N}_{c,s}}{\tilde{N}_{c}}\leq\frac{k^{2}}{\eta sT_{2}}$, we know $R_{i}^{(t)}\leq\tilde{O}(\sigma_{0})$ for $T_{1}+T_{0}<t\leq T_{1}+T_{2}$. ∎